Acids and Bases

1) Which of the following is a general property of a basic solution?

a. feels slippery

b pH less than 7

c. tastes sour

d. turns litmus paper red

e. none of the above

2) If a vinegar sample has a pH of 5, the solution is which of the following:

a. strongly acidic

b. weakly acidic

c neutral

d. weakly basic

e. strongly basic

3) Which of the following is a weak Arrhenius acid

a. HNO3 (aq)

b. HCl (aq)

c HC2H3o2

d. all of the above

e. none of the above

4) If 10.0 mL of M HCl is titrated with 0.200 M NaOH, what volume of sodium hydroxide solution is required to neutralize the acid?

a. 5.00 mL

b. 20.0 mL

c. 40.0 mL

d. 10.0 mL

e. 80.0 mL

5) If a 25.0 mL sample of sulfuric acid is titrated with 50.0mL of 0.200M potassium hydroxide to a phenolphtalein endpoint, what is the molarity of the acid?

a 0.400 M

b. 0.200 M

c. 0.100 M

d. 0.150 M

e. 0.300 M

6) Which of the following is the ionization constant expression for water

a. Kw= (H+)(OH-)

b. Kw = (H2O)/(H+)(OH-)

c. Kw= (H2O) (H20)

d. Kw = (H+) (OH-)(h20)

e. none of the above

7) Which of the following explains why distilled water is neutral?

a distilled water has no ions

b. distilled water does not react with acids

c distilled water has equal concentrations of hydrogen and hydroxide ions

d distilled water has no hydrogen ions

e. none of the above

8) Given an aqueous solution in which the (H+) = 2.0 x 10^-5, what is the molar hydroxide ion concentration?

a (OH-) = 2.0 x 10^-9 M

b. (OH-) = 5.0 x 10^-9 M

c (OH-)= 5.0 x 10^-8 M

d. (OH-) = 2.0 x 10^-10 M

e (OH-) = 5.0 x 10^-10 M
Answer #1

Basic soluhon. eelS ase taste (d) tur. 1.hvur paf, чпе (x) Shova au3) HNO lo Mi020 Molav fMelarl 200 Mz -log 2-0 x1o - S o-30l - 5 (-4-699) 46 ph = 4.699 Poh 14- Ph = 14-4, 699 -30 lo